OpenStudy (anonymous):

Find the exact value: cos (11pi/12) - using half angle identities.

OpenStudy (anonymous):

Half Angle Identity for Cosine: \[Cos(\frac{ a }{ 2 }) = \pm \sqrt{\frac{ 1 + \cos(a) }{ 2 }}\]

OpenStudy (anonymous):

To get it to the form (a/2) you must factor out a half so that your new problem looks like this: \[\cos(\frac{ \frac{ b }{ c } }{ 2 })\] Where B is the new numerator and C is the new denominator and b/c is your "a" in your identity.

OpenStudy (anonymous):

What would be your "a" if you take out a 1/2 from 11pi/12 ?